Mandy buys 5 reams of paper and 3 ink cartridges for $131. Kevin buys 2 reams of paper and 5 ink cartridges for $174. How much does one ream of paper cost? How much does one ink cartridge cost?
Answer by addingup(3677)

5x+3y = 131 (1)
2x+5y = 174 (2)
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Multiply both sides of (1) times 2 and both sides of (2) times 5 and then subtract (2) from (1):
10x+6y = 262
-
10x+25y = 870
-------------
-19y = -608 Divide both sides by -19 and remember that -/- = +
y = 32 is how much ink costs, and:
5x+3(32) = 131
5x+96 = 131 subtract 96 both sides
5x = 35 divide both sides by 5
x = 7 is how much each ream of paper costs
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check: We used equation (1) to find x, so let's try our results on equation (2):
2(7)+5(32) = 174
14+160 = 174 We have the correct answer